OpenStudy (anonymous):

simplify. 35√11-√11 35 34√10 34√11 35√10

OpenStudy (anonymous):

I'm again going to suggest factoring the root. So\[\sqrt{11}(?+1)\]

OpenStudy (gabbyalicorn):

3511−11 =3411 (Decimal: 112.765243)

OpenStudy (gabbyalicorn):

wait...

OpenStudy (anonymous):

oops \[\sqrt{11}(?-1)\]

OpenStudy (gabbyalicorn):

34√11

OpenStudy (anonymous):

----> \[(35-1)\sqrt{11}\]

OpenStudy (anonymous):

Thanks guys